OpenStudy (swaugh):

The perimeter of the triangle is 4x+3y. Find the measure of the missing side. X+y x+y = the two sides

OpenStudy (steve816):

To find the missing side, you will have to subtract the total perimeter by the two of the sides ya?

OpenStudy (steve816):

That would be\[\Large 4x+3y-(x+y+x+y)\]

OpenStudy (swaugh):

Yeah

OpenStudy (swaugh):

I did and I got 2x+1y

OpenStudy (steve816):

is your length of 2 sides x + y + x + y?? You only wrote x + y x+ y

OpenStudy (swaugh):

Am I correct ?

OpenStudy (mathstudent55):

Each side of the other 2 is x + y

OpenStudy (steve816):

Yeah, I just checked your work, you got it right!

OpenStudy (swaugh):

One side = x+y; two side = x+y

OpenStudy (swaugh):

Thanks

OpenStudy (steve816):

You're welcome